\begin{thebibliography}{32}
	\setlength{\itemsep}{0mm} \small
	
	\bibitem{bai2023} W. Bai, W. Liu, Superbiderivations of Simple Modular Lie Superalgebras of Witt Type and Special Type, \emph{Algebra Colloq.} {\bf 30}(02) (2023) 181-192.
	
	\bibitem{Benkovic2009} D. Benkovi\v{c}, Biderivations of triangular algebras, \emph{Linear Algebra Appl.} {\bf 431}(9) (2009) 1587-1602.
	
	\bibitem{Bresar1995} M. Bre\v{s}ar, On generalized biderivations and related maps, \emph{J. Algebra} {\bf 172}(3) (1995) 764-786.
	
	\bibitem{Bresar2018} M. Bre\v{s}ar, K. Zhao, Biderivations and commuting linear maps on Lie algebras, \emph{J. Lie Theory} {\bf 28}(3) (2018) 885-900.
	
	\bibitem{chang20191} Y. Chang, L. Chen, Biderivations and linear commuting maps on the restricted Cartan-type Lie algebras $W(n;\underline{1})$ and $S(n;\underline{1})$, \emph{Linear Multilinear Algebra} {\bf 67}(8) (2019) 1625-1636.

	\bibitem{chang20192} Y. Chang, L. Chen and X. Zhou, Biderivations and linear commuting maps on the restricted Cartan-type Lie algebras $H(n;\underline{1})$, \emph{Comm. Algebra} {\bf 47}(3) (2019) 1311-1326.
	
	\bibitem{chang2021} Y. Chang, L. Chen and Y. Cao, Super-biderivations of the generalized Witt Lie superalgebra $W(m, n;\underline{t})$, \emph{Linear Multilinear Algebra} {\bf 69}(2) (2021) 233-244.
	
	\bibitem{chen2016} Z. Chen, Biderivations and linear commuting maps on simple generalized Witt algebras over a field, \emph{Electron. J. Linear Algebra} {\bf 31}(1) (2016) 1-12.
	
	\bibitem{cheng2017} X. Cheng, M. Wang, J. Sun and H. Zhang, Biderivations and linear commuting maps on the Lie algebra, \emph{Linear Multilinear Algebra} {\bf 65}(12) (2017) 2483-2493.
	
	\bibitem{cheng2019} X. Cheng, J. Sun, Super-biderivations and Linear Commuting Maps on the Twisted $N = 2$ Superconformal Algebra, \emph{Internat. J. Algebra Comput} {\bf 29}(07) (2019) 1235-1247.
	
	\bibitem{Dilxat2023} M. Dilxat, S. Gao and D. Liu. Super-biderivations and Post-Lie Superalgebras on Some Lie Superalgebras, \emph{Acta. Math. Sin. English Ser.} {\bf 39} (2023) 1736-1754. 
	
	\bibitem{du2013} Y. Du, W. Yu, Biderivations of generalized matrix algebras, \emph{Linear Algebra Appl.} {\bf438}(11) (2013) 4483-4499.
	
	\bibitem{fan2017} G. Fan, X. Dai, Super-biderivations of Lie superalgebras, \emph{Linear Multilinear algebra} {\bf 65}(1) (2017) 58-66.
	
	\bibitem{Ghosseiri2013}N.M. Ghosseiri, On biderivations of upper triangular matrix rings, \emph{Linear Algebra Appl.} {\bf 438}(1) (2013) 250-260.
	
	\bibitem{han2016} X. Han, D. Wang and C. Xia, Linear commuting maps and biderivations on the Lie algebras $W (a, b)$, \emph{J. Lie Theory} {\bf 26}(3) (2016) 777-786.
	
	\bibitem{Kac1977} V.G. Kac, Lie superalgebras, \emph{Adv. Math.} {\bf 26} (1977) 8-96.
	
	\bibitem{li2018} W. Li, X. Tang and J. Yuan, Super-biderivations and linear super-commuting maps on the super W-algebra $\widetilde {W}(2, 2) $, \emph{Colloq. Math.} {\bf 153} (2018) 273-300.
	
	\bibitem{Scheunert1979} M. Scheunert, The theory of Lie superalgebras, \emph{Lecture Notes in Math.} {\bf 716} (1979) 3001-454.
	
	\bibitem{tang2017} X. Tang, Biderivations, linear commuting maps and commutative post-Lie algebra structures on W-algebras, \emph{Comm. Algebra} {\bf 45}(12) (2017) 5252-5261.
	
	\bibitem{tang20181}X. Tang, Biderivations and commutative post-Lie algebra structures on the Lie algebra $\mathcal{W}(a,b)$, \emph{Taiwanese J. Math.} {\bf 22}(6) (2018) 1347-1366.

	\bibitem{tang20182} X. Tang, Biderivations of finite-dimensional complex simple Lie algebras, \emph{Linear Multilinear algebra}, {\bf 66}(2) (2018) 250-259.
	
	\bibitem{tang20183} X. Tang,  X. Li, Biderivations of the twisted Heisenberg-Virasoro algebra and their applications, \emph{Comm. Algebra} {\bf 46}(6) (2018) 2346-2355.
	
	\bibitem{tang2020} L. Tang, L. Meng and L. Chen, Super-biderivations and linear super-commuting maps on the Lie superalgebras, \emph{Comm. Algebra} {\bf 48}(12) (2020) 5076-5085.
	
	\bibitem{wang2013} D. Wang, X. Yu, Biderivations and linear commuting maps on the Schr$\ddot{\text{o}}$dinger-Virasoro Lie algebra, \emph{Comm. Algebra} {\bf 41}(6) (2013) 2166-2173.
	
	\bibitem{wang2011} D. Wang, X. Yu and Z. Chen, Biderivations of the parabolic subalgebras of simple Lie algebras, \emph{Comm. Algebra} {\bf 39}(11) (2011) 4097-4104.
	
	\bibitem{xia2016} C. Xia, D. Wang and X. Han, Linear super-commuting maps and super-biderivations on the super-Virasoro algebras, \emph{Comm. Algebra} {\bf 44}(12) (2016) 5342-5350.
	
	\bibitem{xu2015} H. Xu, L. Wang, The properties of biderivations on Heisenberg superalgebras, \emph{Mathematica Aeterna} {\bf 5}(2) (2015) 285-291.
	
	\bibitem{yuan2021} J. Yuan, L. Chen and Y. Cao, Super-Biderivations of Cartan Type Lie Superalgebras, \emph{Comm. Algebra} {\bf 49}(10) (2021) 4416-4426.
	
	\bibitem{yuan2018} J. Yuan, X. Tang, Super-biderivations of classical simple Lie superalgebras, \emph{Aequationes Math.} {\bf 92}(1) (2018) 91-109.
	
	\bibitem{zhang2006} J.H. Zhang, S. Feng, H.X. Li and R.H. Wu, Generalized biderivations of nest algebras, \emph{Linear Algebra Appl.} {\bf 418}(1) (2006) 225-233.
		
	\bibitem{zhang2003} Z.X. Zhang, L.Q. Liu, The centroid of Lie superalgebras, \emph{Journal of Hebei University} {\bf 23}(1) (2003) 85-87.
	
	\bibitem{zhao2020} X. Zhao, Y. Chang, X. Zhou and L. Chen, Super-biderivations of the contact Lie superalgebra $K(m, n;\underline{t})$, \emph{Comm. Algebra} {\bf 488}(8) (2020) 3237-3248.

\end{thebibliography}

\documentclass[8pt,a4paper]{article}
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage[dvips]{graphicx}
\usepackage{latexsym}
\usepackage{mathrsfs}
\usepackage{amscd}
\usepackage{indentfirst}
\usepackage{textcomp}
\usepackage[numbers,square,sort&compress]{natbib} 
\usepackage[colorlinks=true,allcolors=black ]{hyperref}
\allowdisplaybreaks[4]

\newtheorem{Theorem}{Theorem}[section]

\newtheorem{Lemma}[Theorem]{Lemma}

\theoremstyle{definition}
\newtheorem{Definition}{Definition}[section]
\numberwithin{equation}{section}

\newcommand{\BF}{\mathbb{F}}

\begin{document}
\title{{\bf  Super-biderivations and linear super-commuting maps on simple Lie superalgebras}}
\author{\normalsize \bf Da  Xu$^1$\,\,\,\,Qiyuan  Wang$^1$\,\,\,\,Xiaoning  Xu$^1$}
\date{{{\small{  1. School of Mathematics, Liaoning University, Shenyang, 110036,
China  }}}}
\maketitle

\begin{abstract}
Let $L$ be a finite dimensional simple Lie superalgebra over an algebraically closed field of characteristic different from 2. In this paper, we prove that each skew-supersymmtric super-biderivation of $L$ is inner. Furthermore, we prove that each linear super-commuting map on $L$ is a scalar multiplication map.
\end{abstract}
\textbf{Keywords:} Simple Lie superalgebras, Skew-supersymmtric super-biderivations, Linear super-commuting maps.\\
\textbf{2000 Mathematics Subject Classification:} 17B50 17B10

\renewcommand{\thefootnote}{\fnsymbol{footnote}}
\footnote[0]{ Project Supported by National Natural Science
Foundation of China (No.11501274) and the Science Research Project of Liaoning Provincial Education Department, China (No.L2015203).
\\ Author Email: lnuxxn@163.com (X. Xu)}

\section{Introduction}

In recent years, many researchers were interested in biderivations \cite{Benkovic2009,Bresar1995,Bresar2018,chang20191,chang20192,chen2016,cheng2017,du2013,Ghosseiri2013,han2016,tang2017,tang20181,tang20182,tang20183,wang2013,wang2011,zhang2006}. In \cite{Bresar1995}, Bre\v{s}ar showed that every biderivation of right ideals of prime rings is inner and obtained some conclusions about commuting maps and skew-commuting maps. In \cite{Benkovic2009}, Benkovi\v{c} defined the concept of an extremal biderivation, and proved that under certain conditions a biderivation of a triangular algebra is a sum of an extremal and an inner biderivation. Thereafter, Du \cite{du2013} and Wang \cite{wang2011} obtained similar results for generalized matrix algebras and parabolic subalgebras of simple Lie algebras respectively. Since then biderivations of Lie algebras have been studied widely \cite{han2016,tang2017,tang20181,tang20182,tang20183,wang2013}. Furthermore, in \cite{chen2016,cheng2017} the authors proved that each anti-symmetric biderivation is  inner for simple generalized Witt algebras and the Lie algebra $\mathfrak{gca}$. As an application of biderivations, they showed that every linear commuting map is a scalar multiplication map in their respective papers. Afterwards, Chang \cite{chang20191,chang20192} obtained similar results for  restricted Cartan-type Lie algebras $W(n;\underline{1})$, $S(n;\underline{1})$ and $H(n;\underline{1})$.

Lie superalgebras as a generalization of Lie algebras came from supersymmetry in mathematical physics. Naturally, super-biderivations of Lie superalgebras have aroused the interest of many scholars \cite{bai2023,chang2021,cheng2019,Dilxat2023,fan2017,li2018,tang2020,xia2016,xu2015,yuan2021,yuan2018,zhao2020}. In \cite{xu2015}, the definition of super-biderivations of Lie superalgebras  was introduced for the first time and some properties of super-biderivations on Heisenberg superalgebras were described. In \cite{bai2023,chang2021,cheng2019,fan2017,xia2016,zhao2020}, the authors proved that all skew-supersymmetric super-biderivations are inner for simple modular Lie superealgebras of Witt type and special type, generalized Witt Lie superalgebras, twisted $N=2$ superconformal algebras,  centerless super-Virasoro algebras, super-Virasoro algebras and contact Lie superalgebras. Moreover, utilizing the conclusions of super-biderivations, linear super-commuting maps were studied in \cite{bai2023,fan2017,xia2016}. However, by means of the complete classification of classical simple Lie superalgebras over the complex field $\mathbb{C}$, Yuan \cite{yuan2018} proved that all super-biderivations of classical simple Lie superalgebras without the restriction of skew-supersymmetry are inner. Meanwhile, similar results were obtained by them for Cartan type Lie superalgebras over the complex field $\mathbb{C}$ \cite{yuan2021}. As a generalization of the results of Bre\v{s}ar and Zhao on Lie algebras \cite{Bresar2018}, Tang described the intrinsic connections among linear super-commuting maps, super-biderivations and centroids for Lie superalgebras satisfying certain assumptions \cite{tang2020}. 

This paper is devoted to studying the skew-supersymmtric super-biderivations and linear super-commuting maps of a finite dimensional simple Lie superalgebra, which is denoted by $L$, over an algebraically closed field of characteristic different from 2. And the paper is arranged as follows. In Section 2, we review the basic definitions concerning skew-supersymmtric super-biderivations, super-commuting maps and centroids of a Lie superalgebra. In Section 3, we prove that each skew-supersymmtric super-biderivation of $L$ is inner. In Section 4, we prove that each linear super-commuting map on $L$ is a scalar multiplication map.

\section{Preliminaries}

In this section, we recall some basic notations concerning Lie superalgebras \cite{Scheunert1979,Kac1977}.

Suppose $\mathbb{F}$ is an algebraically closed field of characteristic different from 2. Let  $L=L_{\bar{0}} \oplus L_{\bar{1}}$ be a Lie superalgebra over $\mathbb{F}$. A linear map $f : L \rightarrow L$ is said to be homogeneous of degree $\gamma$, $\gamma\in\mathbb{Z}_2$, if
  \[f(L_\alpha)\subset L_{\alpha+\gamma},\]
for any $\alpha\in\mathbb{Z}_2$. And we write $|f|=\gamma$  for the degree of the homogeneous linear map $f$. If $\gamma=0$, then $f$ is called an even linear map. Throughout this paper, we should mention that once the symbol $|f|$ or $|x|$ appears in an expression, it implies that $f$  is a homogeneous linear map or $x$ is a homogeneous element. A linear map $D : L \rightarrow L$ is called a superderivation of $L$, if
  \[D([x,y])=[D(x),y]+(-1)^{|D||x|}[x,D(y)],\]
for any $x,y\in L$. A bilinear mapping $b : L\times L \rightarrow L$ is said to be homogeneous of degree $\gamma$, $\gamma\in\mathbb{Z}_2$, if
  \[b(L_\alpha,L_\beta)\subset L_{\alpha+\beta+\gamma},\]
for any $\alpha,\beta\in\mathbb{Z}_2$. Similarly, we write $|b|=\gamma$  for the degree of the homogeneous bilinear map $b$. If $\gamma=0$, then $b$ is called an even bilinear map. A bilinear map $\phi : L\times L \rightarrow L$ is called a skew-supersymmtric super-biderivation of $L$, if
  \begin{align*}
  	&\phi(x,y)=-(-1)^{|x||y|}\phi(y,x),\\
  	&\phi(x,[y,z])=[\phi(x,y),z]+(-1)^{(|\phi|+|x|)|y|}[y,\phi(x,z)], 
  \end{align*}
for any $x,y,z\in L$. 	If the bilinear map $\phi_{\lambda}$ : $L\times L\rightarrow L$ is defined by
  \[\phi_{\lambda}( x,y ) =\lambda [ x,y],\]
for any $x,y\in L$, where $\lambda \in \BF$. Obviously $\phi_{\lambda}$ is a skew-supersymmtric super-biderivation of $L$. This class of super-biderivations is called inner super-biderivations. 

\begin{Definition}\cite{tang2020}
	For a Lie superalgebra $L=L_{\bar{0}} \oplus L_{\bar{1}}$, an even linear map $f : L \rightarrow L$ is called a linear super-commuting map if
	  \[[f(x),y]=[x,f(y)],\]
	for any $x,y\in L$.
\end{Definition}

\begin{Definition}\cite{zhang2003}\label{def2.2}
	The centroid of a Lie  superalgebra $L=L_{\bar{0}} \oplus L_{\bar{1}}$ is the associative superalgebra
	  \[\Gamma(L)=\{f\in pl(L)\mid [f,adx]=0,\ \text{for any}\ x\in L\}.\]
	Observe that $\Gamma(L)$ is a associative superalgebra over the field $\mathbb{F}$. Then
	   \[\Gamma(L)=\Gamma(L)_{\bar{0}} \oplus \Gamma(L)_{\bar{1}},\] 
	 where $\Gamma(L)_\alpha=\Gamma(L)\bigcap pl_\alpha(L),\alpha\in\mathbb{Z}_2$.
\end{Definition}

\section{Skew-supersymmtric super-biderivations}

 In this section, we establish several lemmas about the skew-supersymmtric super-biderivations of the simple Lie superalgebra $L$. Moreover, we prove that each skew-supersymmtric super-biderivation of $L$ is inner.
 
 \begin{Lemma}\label{lem3.1}
 	Suppose that L is a finite dimensional simple Lie superalgebra over $\mathbb{F}$. Then 
 	  \[\Gamma(L)_{\bar{0}}=\mathbb{F}\cdot \mathrm{Id},\ \Gamma(L)_{\bar{1}}=0,\]
 	where $\mathrm{Id}$ is the identity map on $L$. And further we have $\Gamma(L)=\mathbb{F}\cdot \mathrm{Id}$.
 	\begin{proof}
 		By the Definition \ref{def2.2}, for any $f\in \Gamma(L)$, we have that
 		  \[f\circ adx=(-1)^{|f||x|}adx\circ f,\]
 		for any $x\in L$. Above equation is equivalent to 
 		  \[f([x,y])=(-1)^{|f||x|}[x,f(y)]=[f(x),y],\]
 		for any $x,y\in L$. 
 		
 		(I) Obviously $\Gamma(L)_{\bar{0}}\supset\mathbb{F}\cdot \mathrm{Id}$, we only need to prove $\Gamma(L)_{\bar{0}}\subset\mathbb{F}\cdot \mathrm{Id}$. Set $0\neq f\in \Gamma(L)_{\bar{0}}$. Due to $\mathbb{F}$ is an algebraically closed field, $f$ has a nonzero eigenvalue $\lambda\in\mathbb{F}$. Let $E_\lambda$ denote the corresponding eigen-subspace. For any $x\in L,y\in E_\lambda$, we hvae
 		\begin{align*}
 			f([x,y])=[x,f(y)]=[x,\lambda y]=\lambda[x,y];\\
 			f([y,x])=[f(y),x]=[\lambda y,x]=\lambda[y,x].
 		\end{align*}
 	    According to the above two equations, it follows that $[L,E_\lambda]\subset E_\lambda$ i.e. $E_\lambda$ is an ideal of $L$. Due to $L$ is simple and $E_\lambda\neq0$, we get that $E_\lambda=L$ i.e. $f=\lambda\cdot\mathrm{Id}\in\mathbb{F}\cdot \mathrm{Id}$.
 	  
 	    (II) Set $f\in \Gamma(L)_{\bar{1}}$. For any $x,y\in L$, we have
 	      \[f([x,y])=(-1)^{|x|}[x,f(y)]=[f(x),y].\]
 	    It is easy to see that the kernel of $f$ is an ideal of $L$, hence either $f= 0$ or else $f$ is bijective. On the one hand,
 	      \[f\circ f([x,y])=(-1)^{|x|}f([x,f(y)])=(-1)^{|x|}[f(x),f(y)].\]
 	    On the other hand,
 	      \[f\circ f([x,y])=f([f(x),y])=(-1)^{\bar{1}+|x|}[f(x),f(y)].\]
 	    Comparing two sides of the above two equations, we have 
 	      \[f\circ f([x,y])=-f\circ f([x,y]),\]
 	    for any $x,y\in L$. Due to $L$ is simple and the characteristic of $\mathbb{F}$ is different from 2, we get 
 	      \[f\circ f=0.\]
 	    Thus we have shown that $f=0$.
 	\end{proof}
 \end{Lemma}

\begin{Lemma}\label{lem3.2}
	Suppose that L is a simple Lie superalgebra over $\mathbb{F}$. Then every skew-supersymmtric superbiderivation $\phi : L\times L \rightarrow L$ can be written as
	  \[\phi(x,y)=f_\phi([x,y]),\]
	where $f_\phi\in \Gamma(L)$.
	\begin{proof}
		Because $L$ is a centerless Lie superalgebra and $L=[L,L]$. Then the conclusion holds from Theorem 3.8 in \cite{tang2020}. 
	\end{proof}
\end{Lemma}

\begin{Theorem}\label{the3.3}
	Suppose that L is a finite dimensional simple Lie superalgebra over $\mathbb{F}$. Then each skew-supersymmtric superbiderivation of $L$ is inner.
	\begin{proof}
		Suppose $\phi : L\times L \rightarrow L$ is a skew-supersymmtric superbiderivation of $L$. By Lemma \ref{lem3.1} and Lemma \ref{lem3.2}, for any $x,y\in L$, there is an element $\lambda\in\mathbb{F}$ such that
		  \[\phi(x,y)=f_\phi([x,y])=\lambda[x,y],\]
		where  $\lambda$ is dependent on $\phi$.
	\end{proof}
\end{Theorem}

\section{Linear super-commuting maps}

In this section, by using the conclusion of skew-supersymmtric superbiderivations we prove that each linear super-commuting map on $L$ is a scalar multiplication map.

\begin{Lemma}\label{lem4.1}
	Suppose that L is a Lie superalgebra over $\mathbb{F}$. Let  $f : L \rightarrow L$ be a linear super-commuting map and 
	  \[\phi_f(x,y)=[f(x),y],\]
	for any $x,y\in L$. Then $\phi_f : L\times L \rightarrow L$ is a skew-supersymmtric super-biderivation.
	\begin{proof}
		It is easy to see that $|\phi_f |=|f|=\bar{0}$. For one thing, we get
		  \begin{align*}
		  	\phi_f (x,y)&=[f(x),y]=[x,f(y)]=-(-1)^{|x||y|}[f(y),x]\\
		  	              &=-(-1)^{|x||y|}\phi_f (y,x),
		  \end{align*}
	    for any $x,y\in L$. For another, we have
	      \begin{align*}
	      	\phi_f (x,[y,z])&=[f(x),[y,z]]\\
	      	                   &=[[f(x),y],z]+(-1)^{(|f|+|x|)|y|}[y,[f(x),z]]\\
	      	                   &=[\phi_f (x,y),z]+(-1)^{(|\phi_f |+|x|)|y|}[y,\phi_f (x,z)].
	      \end{align*}
        for any $x,y,z\in L$.
	\end{proof}
\end{Lemma}

\begin{Theorem}
	Suppose that L is a finite dimensional simple Lie superalgebra over $\mathbb{F}$. Then each linear super-commuting map on $L$ is a scalar multiplication map.
	\begin{proof}
		Let  $f : L \rightarrow L$ be a linear super-commuting map. By Theorem \ref{the3.3} and Lemma \ref{lem4.1}, for any $x,y\in L$, there ia an element $\lambda\in\mathbb{F}$ such that
		  \[[f(x),y]=\phi_f (x,y)=\lambda[x,y],\]
		where  $\lambda$ is dependent on $f$. This implies that $[f(x)-\lambda x, y]=0$ for any $y\in L$. Due to $L$ is simple, it is easy to see that
		  \[f(x)=\lambda x,\]
		for any $x\in L$.
	\end{proof}
\end{Theorem}

%

\end{document}